Solve the following equation :
(1) (D^2 - 2DD' + [D']^2 )z = tan(x - y)
(2) z(p - q) = z^2 + (x + y^2)
1
Expert's answer
2020-03-20T17:29:57-0400

Question (i)

"\left(D^2-2DD'+[D']^2\right)z=\tan(x-y)\longrightarrow\\[0.3cm]
z=C.F.+P.I."



1 STEP: We try to find C.F.



"\left(D^2-2DD'+[D']^2\right)z=0"



The auxillary equation is


"m^2-2m+1=0\longrightarrow\\[0.3cm]
m_{1,2}=1"



Then,


"C.F.=f_1(y+1\cdot x)+x\cdot f_2(y+1\cdot x)\\[0.3cm]
\boxed{C.F.=f_1(y+x)+x\cdot f_2(y+x)}"



2 STEP: We try to find P.I.

Let us give some theory to understand how to solve this equation:

If "F(x,y)"  is any function, resolve "F(D,D')" into linear factors say



"(D-m_1D'),(D-m_2D'),\ldots,(D-m_nD')"



then,



"P.I.=\displaystyle\frac{1}{(D-m_1D')\cdot\ldots(D-m_nD')}F(x,y)"



Note:



"(1)\,\,\frac{1}{(D-mD')}F(x,y)=\int F(x,c-mx)dx,\,\,\text{where}\,\,y=c-mx\\[0.3cm]
(1)\,\,\frac{1}{(D+mD')}F(x,y)=\int F(x,c+mx)dx,\,\,\text{where}\,\,y=c+mx\\[0.3cm]"



In our case,



"P.I.=\frac{1}{(D-1\cdot D')(D-1\cdot D')}\tan(x-y)=\\[0.3cm]
=\frac{1}{(D-1\cdot D')}\int\tan(x-(c-1x))dx=[1x+y=c]=\\[0.3cm]
=\frac{1}{(D-1\cdot D')}\int\tan(2x-c)dx=\\[0.3cm]
=\frac{1}{(D-1\cdot D')}\left(\frac{-1}{2}\cdot\ln|\cos(2x-c)|\right)=\\[0.3cm]
=\int\left(\frac{-1}{2}\cdot\ln|\cos(2x-c)|\right)dx"



(More information about "\int\tan(x)dx": https://en.wikipedia.org/wiki/List_of_integrals_of_trigonometric_functions)

Conclusion,



"\boxed{P.I.=-\frac{1}{2}\cdot\int\ln|\cos(2x-c)|dx}"



General conclusion,



"z=C.F.+P.I.\longrightarrow\\[0.3cm]
\boxed{z=f_1(y+x)+x\cdot f_2(y+x)-\frac{1}{2}\cdot\int\ln|\cos(2x-c)|dx}"



Question (ii)



"z(p-q)=z^2+(x+y)^2\longrightarrow\\[0.3cm]
zp+(-z)q=(z^2+(x+y)^2)"



The auxillary equation is



"\frac{dx}{z}=\frac{dy}{-z}=\frac{dz}{z^2+(x+y)^2}\\[0.3cm]"

Then,



"\frac{dx}{z}=\frac{dy}{-z}\longrightarrow\frac{dx}{1}=\frac{dy}{-1}\longrightarrow\\[0.3cm]
dx+dy=0\longrightarrow d(x+y)=0\longrightarrow\\[0.3cm]
\boxed{x+y=C_1}\\[0.3cm]
\frac{dx}{z}=\frac{dz}{z^2+(x+y)^2}\longrightarrow dx=\frac{zdz}{z^2+C_1^2}\longrightarrow\\[0.3cm]
x=\frac{1}{2}\cdot\int\frac{d(z^2)}{z^2+C_1^2}\longrightarrow x=\frac{1}{2}\cdot\ln|z^2+C_1^2|-\frac{1}{2}\ln|C_2|\longrightarrow\\[0.3cm]
2x=\ln\left|\frac{z^2+C_1^2}{C_2}\right|\longrightarrow e^{2x}=\frac{z^2+C_1^2}{C_2}\\[0.3cm]
\boxed{C_2=\frac{z^2+C_1^2}{e^{2x}}}\\[0.3cm]"

Conclusion,



"\boxed{\text{Solution of equation is}\,\,\varphi\left((x+y),e^{-2x}(z^2+(x+y)^2)\right)=0}"



ANSWER

Question (i)



"z=f_1(y+x)+x\cdot f_2(y+x)-\frac{1}{2}\cdot\int\ln|\cos(2x-c)|dx"

Question (ii)



"\varphi\left((x+y),e^{-2x}(z^2+(x+y)^2)\right)=0"